# Galvanometer-scanning transient phase microscopy with balanced detection and arbitrary pump polarization

### v1

@@ 3.1 Transient Phase and Absorption Measurements
This section presents the results of measuring transient absorption and transient phase in glass, a graphene monolayer, bovine hemoglobin in solution and in live human red blood cells. The probe wavelength of 800 nm is considered off-resonance for the hemoglobin and red blood cell samples.
@@ 3.1.1 Cross-Phase Modulation Measurement in Glass
First we acquired a delay scan in a glass microscope slide (Figure 5). Glass has a minimal absorption at the probe wavelength of 800 nm and the interaction that occurs between the sample and the pump pulse is instantaneous. Thus, the width of the XPM pulses gives us an estimate of pulse duration for our reference and probe pulses as they overlap with the pump pulse in time; these pulses are 350 fs in duration.
@@ 3.1.2 Graphene Monolayer
Next we explored the 2D and 3D nonlinear response of a graphene monolayer on a 0.5 mm quartz substrate placed on top of a 1 mm glass microscope slide. Previous works looking at graphene have utilized transient absorption to analyze graphene carrier dynamics [GrapheneTAS_SingleLayerHyperspectral_Gesuele_2019, MaterialSciApp_3_Ma_2022]. It is a material that shows promise for use in various applications including electronics and drug delivery, making it a popular research target [Graphene_TAM_InBlood_Yang_2015]. Here, graphene serves as a control material to compare TAM with TM since it has a strong transient absorption response at our pump and probe wavelengths.
@@
An initial delay scan taken in the center of the sample revealed a stronger transient absorption versus transient phase response with a decay time of 2 picoseconds.
The first peak at ps corresponds to the pump-reference overlap and the second peak at ps corresponds to the pump-probe overlap.
@@
Image stacks were acquired at the edge of the graphene sample where we expected to see some variation in the deposition of graphene on the quartz substrate. A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e. ) with the pump blocked. From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared with output from the lock-in amplifier (V) (Figure 8). The transient absorption measurements still provide better SNR, but similar structural information can be seen in both techniques.
@@ 3.1.3 Hemoglobin and Red Blood Cells
Next we moved onto biological samples. First, we looked at the spectroscopy of a hemoglobin sample prepared with 9 mM of bovine hemoglobin lyophilized powder (Sigma-Aldrich H2500) in phosphate buffer solution. Then we imaged a droplet of whole blood placed on a poly-L-lysine coated microscope slide with a glass coverslip. The whole blood used was acquired from healthy donors by venipuncture by Dr. Liszt Madruga of the Department of Chemical Engineering (CSU) working under Prof. Matt Kipper of the Departments of Chemical & Biological Engineering (CSU) and Prof. Ketul Popat of the Department of Mechanical Engineering (CSU). The protocol followed was approved by the Colorado State University Institutional Review Board.
@@
The hemoglobin sample showed a decay curve (top row in Figure 9). Interestingly, hemoglobin demonstrated a pump polarization-dependent signal with differences between the polarization sensitivity of the transient absorption versus transient phase measurements (bottom row of Figure 9). This would be impossible to observe if the pump-combining dichroic was placed before the probe-reference splitting calcite like in [Orrit, Guillet_2015]. With transient absorption the pump-reference overlap showed a polarization dependence while the pump-probe overlap did not. For the pump-reference overlap, the strongest signal occurred with an s-polarized pump. With transient phase both overlaps demonstrated a polarization dependence; the pump-reference overlap shows the largest signal for the s-polarized pump and the pump-probe overlap shows the largest signal for the p-polarized pump.
@@
Imaging human red blood cells, we expected to see a predominant signal from hemoglobin and a similar SNR from the two techniques (based on the results from the delay scans). However, the transient phase demonstrated a stronger signal than transient absorption in this case with clearer structural details due to XPM. Figure 10 shows a comparison of TM and TAM for imaging the transient decay in red blood cells. All of the images are shown on the same intensity scale; the overlaps with the pump-reference at ps and with the pump-probe at ps show the strongest signals and good structural detail for TM.
@@ 4.1 Impacts of Dichroic Beam-Combiner on SNR
The pump-combining dichroic was placed after the probe-reference splitting calcite, unlike prior works which placed the dichroic before the splitter to avoid splitting the pump [Orrit, Guillet_2015]. This enabled characterization of polarization dependence, as well as selection of relative pump-probe polarization that maximizes signal for imaging. It should be noted, however, that the dielectric coating of the dichroic is subject to polarization aberrations, including dichroism (polarization-dependent absorption), depolarization, dispersion, and retardance [DichroicAberrations]. These effects can depend on both incident polarization state and angle of incidence, and can ultimately interfere with our detection scheme. One behavior we observed with our setup was a polarization-dependent spectral phase difference imparted to the probe and reference from the dichroic; the pulse broadening effects were minimal and inconsequential with our pulse durations (>300 fs), though we did notice a variation in spectral fringe visibility as a function of wavelength and the development of ellipticity on the probe and reference pulses. Additionally, our reference and probe pulses were found to have unequal amplitudes at the detectors. For TAM, only the amplitude differences contribute to a reduced SNR due to a DC offset. For TM, all three effects contribute to a reduced SNR. Thus, future work to address these issues will likely further the advantage of TM over TAM in this scenario.
@@ 4.2 Impact of Galvanometer-scanning on Re-timing Phase in Imaging
Galvanometer-scanning microscopy is a new development in TM and comes at the price of an angle-dependent re-timing phase present across the scanned field of view. In order to determine how great an impact galvanometer-dependent re-timing phase has on this system, a cross-phase modulation (XPM) image was taken in a glass slide at different fields of view. Since glass is homogeneous and has no transient signal, a flat signal would be expected across the entire field of view in the absence of galvanometer-dependent re-timing phase. The measured data shown in Figure 11 indicates a phase shift that depends on scan angle. The images shown in Figure 11 have had the on-axis phase shift DC subtracted so only the spatially dependent phase shift is portrayed.
Clearly, the larger field of view sees a stronger galvanometer-dependent phase shift . Overall, the signal measured across a field of view will equal the sum + . Decreasing the field of view or increasing the magnification of the beam at the re-timing calcite (CAL 2) could help minimize the re-timing phase effects (Supplementary 4). In order to completely eliminate this effect, de-scanning could be added in prior to CAL 2.
@@ 4.3 Trade-offs
One significant benefit of measuring TM is that the instantaneous cross-phase modulation (XPM) is measured in addition to the transient phase signal. The XPM contributes additional context about the structure of the sample related to nonlinear refractive index variations across the sample [XPMSS_wilson]. TAM will also measure instantaneous phenomena, such as two-photon absorption, but this does not give the same structural clarity as XPM and is instead specific only to materials and molecules resonant with the sum of the pump and probe energies.
@@
For thin, weakly birefringent samples the polarization relationship of the reference and probe pulses is at risk of changing, and we would expect the effects to be similar to the dichroic in reducing SNR. Thick birefringent samples run the risk of not only changing the polarization relationship of the reference and probe pulses but also of additional pulse splitting or temporal separation. Depending on the degree to which this occurs, a very small amount of the exiting reference and probe might be within the re-timing range of the secondary calcite, or the pulses may not be able to be re-timed by the calcite at all. This would preclude the ability of this system to make phase measurements entirely.
@@ 5 Conclusion
The purpose of this research was to extend TM to a galvanometer-scanning configuration, as a complement to TAM for biological samples. In sum, the scanning galvanometer TM approach described here can offer advantages over TAM, such as XPM structural contrast and potentially higher signal levels, but is susceptible to degradation from polarization aberrations in the optical setup and sample birefringence, and has more limited field of view. Thus, it is beneficial to be able to acquire both TAM and TM with a simple rotation of a waveplate, as described here. In the future two changes to the microscope design here could be made to improve the utility for TM. Firstly, the dichroic beam combiner could be swapped for a 50/50 beamsplitter. While this would cut the power at the combiner by half, it would also eliminate the polarization-dependent effects imparted by the dichroic that we have shown negatively impact the SNR of TM. Secondly, de-scanning could be added prior to the re-timing calcite to remove the spatial dependence on re-timing phase. Overall, TM shows promise as a complementary technique to TAM, particularly for setups that may be limited in their ability to vary probe wavelengths.

### v2

@@ 3.1 Transient Phase and Absorption Measurements
This section presents the results of measuring transient absorption and transient phase in glass, a graphene monolayer, bovine hemoglobin in solution and fresh human red blood cells.
@@ 3.1.1 Cross-Phase Modulation Measurement in Glass
First we acquired a delay scan in a glass microscope slide (Figure 5). Glass has a minimal absorption at the probe wavelength of 800 nm and the interaction that occurs between the sample and the pump pulse is instantaneous. When making an absorption measurement, we would expect to see a two-photon absorption peak when the pump and probe pulses were overlapped if the sample absorbed the pump wavelength. For a phase measurement, we are able to see cross-phase modulation (XPM) from the instantaneous optical Kerr effect, in which pump intensity changes the refractive index seen by the probe [21]. Consequently, due to the instantaneous nature of the interaction, the width of the XPM pulses indicates a cross-correlation between pump and probe. (Frequency resolved optical gating (FROG) measurements were previously made that showed the probe pulse width was the limiting factor).
@@ 3.1.2 Graphene Monolayer
Next we examined a graphene monolayer on a 0.5 mm quartz substrate placed on top of a 1 mm glass microscope slide. Previous works looking at graphene have utilized transient absorption to analyze graphene carrier dynamics [39, 12]. Near-infrared pumping (as in our setup) results in transient heating of free carriers [40]. It is a material that shows promise for use in various applications including electronics and drug delivery, making it a popular research target [41].
@@
Figure 6 shows a delay scan taken in the center of the sample. A stronger transient absorption versus transient phase response is observed, with a decay time of 2 ps (found by fitting to an exponential function). The first peak at ps corresponds to the pump-reference overlap and the second peak at ps corresponds to the pump-probe overlap.
@@
Image stacks were acquired at the edge of the graphene sample where spatially-varying structure is expected. Images of pixels were acquired with 0.05 second pixel dwell time (8.3 minutes/frame). A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e. ) with the pump blocked. From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared (Fig. 8). Transient absorption still provides better SNR, but similar structural information can be seen in both.
@@ 3.1.3 Hemoglobin and Red Blood Cells
Next we moved onto biological samples. First, we looked at the spectroscopy of a hemoglobin sample prepared with 9 mM of bovine hemoglobin lyophilized powder (Sigma-Aldrich H2500) in phosphate buffer solution.
Second, we imaged a droplet of whole blood placed on a poly-L-lysine coated microscope slide with a glass coverslip. The whole blood used was acquired from healthy donors by venipuncture by Dr. Liszt Madruga of the Department of Chemical Engineering (CSU) working under Prof. Matt Kipper of the Departments of Chemical & Biological Engineering (CSU) and Prof. Ketul Popat of the Department of Mechanical Engineering (CSU). The protocol followed was approved by the Colorado State University Institutional Review Board.
@@
The hemoglobin sample showed a decay curve (top row, Fig. 9). For a discussion of the physics behind the transient absorption response of hemoglobin at these wavelengths, we refer the reader to [6]. Interestingly, hemoglobin demonstrated a pump polarization-dependent signal with differences between the polarization sensitivity of the transient absorption versus transient phase measurements (bottom row, Fig. 9). This would be impossible to observe if the pump-combining dichroic was placed before the probe-reference splitting calcite like in [32, 33]. With transient absorption the pump-reference overlap showed a polarization dependence while the pump-probe overlap did not. For the pump-reference overlap, the strongest signal occurred with an s-polarized pump. With transient phase both overlaps demonstrated a polarization dependence; the pump-reference overlap shows the largest signal for the s-polarized pump and the pump-probe overlap shows the largest signal for the p-polarized pump. The polarization dependence is stronger at shorter delays (< 1 ps), similar to pump-probe anisotropy measurements on other hemeproteins [42].
@@
Imaging human red blood cells, we expected to see a predominant signal from hemoglobin and a similar signal level from the two techniques (based on the results from the delay scans). However, TM demonstrated a much larger signal, with clearer structural details (Fig. 10). All of the images are shown on the same intensity scale; the overlaps with the pump-reference at ps and with the pump-probe at ps show the strongest signals and good structural detail for TM.
@@ 4.1 Impacts of Dichroic Beam-Combiner on SNR
The pump-combining dichroic was placed after the probe-reference splitting calcite, unlike prior works which placed the dichroic before the splitter to avoid splitting the pump [32, 33]. This enabled characterization of polarization dependence, as well as selection of relative pump-probe polarization that maximizes signal for imaging. It should be noted, however, that the dielectric coating of the dichroic is subject to polarization aberrations, including dichroism (polarization-dependent absorption), depolarization, dispersion, and retardance [37]. These effects can depend on both incident polarization state and angle of incidence, and can ultimately interfere with our detection scheme. One behavior we observed with our setup was a polarization-dependent spectral phase difference imparted to the probe and reference from the dichroic; the pulse broadening effects were minimal and inconsequential with our pulse durations (>300 fs), though we did notice a variation in spectral fringe visibility as a function of wavelength and the development of ellipticity on the probe and reference pulses. Additionally, our reference and probe pulses were found to have unequal amplitudes at the detectors. For TAM, only the amplitude differences contribute to a reduced SNR due to a DC offset. For TM, all three effects contribute to a reduced SNR. Thus, future work to address these issues will likely further the advantage of TM over TAM in this scenario.
@@ 4.2 Impact of Galvanometer-scanning on Re-timing Phase in Imaging
One drawback to this TM approach is an angle-dependent re-timing phase offset across the scanned field of view (FOV). To characterize this, an XPM image was taken in a glass slide at different FOV. Since glass is homogeneous and has no transient signal, a flat signal would be expected across the entire FOV in the absence of spatially-dependent re-timing. The measured data shown in Figure 11 indicates a phase shift that depends on scan angle. The images shown in Figure 11 have had the on-axis (center of image) phase shift subtracted so only the spatially dependent phase shift is portrayed.
Clearly, the larger FOV sees a stronger galvanometer-dependent phase shift . Overall, the signal measured across a FOV will equal the sum + . Decreasing the FOV or increasing the magnification of the beam at the re-timing calcite (CAL 2) could help minimize the re-timing phase effects (Supplementary 4). In order to completely eliminate this effect, de-scanning could be added in prior to CAL 2.
@@ 4.3 Trade-offs
One significant benefit of measuring TM is that XPM is measured in addition to pump-induced transients. The XPM contributes additional context about the structure of the sample related to nonlinear refractive index variations across the sample [29], though it carries no information about excited-state lifetimes. TAM will also measure instantaneous phenomena, such as two-photon absorption, but this does not give the same structural clarity as XPM and is instead specific only to materials and molecules resonant with the sum of the pump and probe energies (with peak intensities sufficient for exciting nonlinear processes).
@@
Another potential drawback to this setup is the need for polarization purity between the probe and reference. Any loss in the ability to separate probe and reference could degrade SNR, e.g. from the dichroic, microscope optics, and the sample. Thick birefringent samples run the risk of not only changing the polarization relationship of the reference and probe pulses but also of additional pulse splitting or temporal separation. Depending on the degree to which this occurs, a very small amount of the exiting reference and probe might be within the re-timing range of the secondary calcite, or the pulses may not be able to be re-timed by the calcite at all. This would preclude the ability of this system to make phase measurements entirely.
@@ 5 Conclusion
The purpose of this research was to extend TM, a measurement of pump-induced refractive index changes measured with interferometric probe and reference pulses, to a galvanometer-scanning configuration as a complement to TAM for biological samples. In sum, TM can offer advantages over TAM, such as XPM structural contrast and potentially higher signal levels (as was the case here for hemoglobin and red blood cells), but is susceptible to degradation from polarization aberrations in the optical setup and sample birefringence, and has more limited FOV. Thus, it is beneficial to be able to acquire both TAM and TM with a simple rotation of a waveplate, as described here. In the future, two changes could be made to improve the utility for TM. Firstly, the dichroic beam combiner could be swapped for a 50/50 beamsplitter. While this would cut the power at the combiner by half, it would also eliminate the polarization-dependent effects imparted by the dichroic that we have shown negatively impact the SNR of TM. Secondly, de-scanning could be added prior to the re-timing calcite to remove the spatial dependence on re-timing phase. Overall, TM shows promise as a complementary technique to TAM, particularly for setups that may be limited in their ability to vary probe wavelengths and as a way to enhance structural context of a transient image through its simultaneous measurement of XPM.